# FaceXFormer: A Unified Transformer for Facial Analysis

### v2

@@ Abstract
In this work, we introduce FaceXFormer, an end-to-end unified transformer model capable of performing nine facial analysis tasks including face parsing, landmark detection, head pose estimation, attribute prediction, and estimation of age, gender, race, expression, and face visibility within a single framework. Conventional methods in face analysis have often relied on task-specific designs and pre-processing techniques, which limit their scalability and integration into a unified architecture. Unlike these conventional methods, FaceXFormer leverages a transformer-based encoder-decoder architecture where each task is treated as a learnable token, enabling the seamless integration and simultaneous processing of multiple tasks within a single framework. Moreover, we propose a novel parameter-efficient decoder, FaceX, which jointly processes face and task tokens, thereby learning generalized and robust face representations across different tasks. We jointly trained FaceXFormer on nine face perception datasets and conducted experiments against specialized and multi-task models in both intra-dataset and cross-dataset evaluations across multiple benchmarks, showcasing state-of-the-art or competitive performance. Further, we performed a comprehensive analysis of different backbones for unified face task processing and evaluated our model “in-the-wild”, demonstrating its robustness and generalizability. To the best of our knowledge, this is the first work to propose a single model capable of handling nine facial analysis tasks while maintaining real-time performance at $33.21$ FPS.
Figure 1: FaceXFormer an end-to-end unified transformer model for 9 different facial analysis tasks such as face parsing, landmark detection, head pose estimation, attributes recognition, and estimation of age, gender, race, expression and face visibility.
@@ 1 Introduction
Face analysis is a crucial problem as it has broad range of application such as face verification and identification [84, 85], surveillance [22], face swapping [13], face editing [127], de-occlusion [111], 3D face reconstruction [102], retail [1], image generation [109] and face retrieval [113]. Facial analysis tasks (Figure 1 involve face parsing [28, 98], landmarks detection [46, 125], head pose estimation [124, 12], facial attributes recognition [64, 57], age/gender/race/expression estimation [8, 43] and landmarks visibility prediction [54, 35]. Therefore, developing a generalized and robust face model for all tasks is a crucial and longstanding problem in the face community.
In recent years, significant advancements have been made in facial analysis, developing state-of-the-art methods and face libraries for various tasks  [124, 125, 43, 12, 111, 13]. Despite these methods achieving promising performance, they cannot be integrated into a single pipeline due to their specialized model designs and task-specific pre-processing techniques. Furthermore, deploying multiple specialized models simultaneously is computationally intensive and impractical for real-time applications, leading to increased system complexity and resource consumption. These challenges emphasis the need for a unified model that can concurrently handle multiple facial analysis tasks efficiently (see Table 1). A single model capable of addressing multiple facial tasks is desirable because it: (1) learns a robust and generalized face representation capable of handling in-the-wild images; (2) intra-task modeling helps the models to learn task-invariant representation; and (3) simplifies deployment pipelines by reducing computational overhead and achieving real-time performance.
Methods FP LD HPE Attr Age Gen Race Vis Exp
@@
Table 1: Comparison with representative methods under different task settings. Our proposed FaceXFormer can perform various facial analysis tasks in a single model. FP - Face Parsing, LD - Landmarks Detection, HPE - Head Pose Estimation, Attr - Attributes Recognition, Age - Age, Gen - Gender, Race - Race Estimation, Exp - Facial Expression Recognition, and Vis - Face Visibility
To this end, we introduce FaceXFormer, an end-to-end unified model designed for nine different facial analysis tasks, such as face parsing, landmark detection, head pose estimation, attribute recognition, age/gender/race/expression estimation and face visibility prediction. FaceXFormer enables task unification by leveraging the transformers and learnable tokens as its core components. Specifically, we introduce a transformer-based encoder-decoder structure, treating each facial analysis task as a unique, learnable token within the framework. Treating each task as a token allows for the simultaneous processing of multiple facial analysis tasks, overcoming the challenges present in conventional methods that depend on separate, task-specific models and pre-processing routines. Furthermore, we introduce a parameter-efficient decoder, FaceX, which processes both face and task tokens together, enabling the model to learn robust face representations that generalize across various tasks. This parameter-efficient design reduces computational load and allows our model to perform in real-time. After modeling the intra-task and face-token relationships in the FaceX decoder, the task tokens are fed into a unified head, which essentially converts these task tokens into corresponding task predictions.
Our extensive experiments demonstrate that FaceXFormer achieves state-of-the-art or competitive performance compared to specialized models and existing multi-task frameworks, including both intra-dataset and cross-dataset evaluations across multiple benchmarks. Specifically, recent multi-task frameworks such as Faceptor [67], QFace [83], and Swinface [66] address fewer tasks than FaceXFormer, yet our model outperforms them on existing benchmarks. Moreover, we show that our model effectively handles images ’in the wild,’ demonstrating its robustness and generalizability across nine different tasks. This robustness is critical for real-world applications where uncontrolled conditions and diverse inputs are common. Achieving real-time performance at $33.21$ FPS marks a significant advancement, making FaceXFormer highly suitable for time-sensitive applications. To the best of our knowledge, this is the first work to propose a single model capable of handling nine different facial analysis tasks using transformers, all while maintaining real-time performance.
@@
    We introduce FaceXFormer, a unified transformer-based framework capable of simultaneously processing nine different facial analysis tasks, achieving real-time performance of $33.21$ FPS.
@@
    We propose FaceX, a parameter-efficient decoder that represents each facial analysis task as a token, enabling joint processing of face and task tokens.
@@
    We conduct extensive experiments and analyses, including both intra-dataset and cross-dataset evaluations, demonstrating that our approach achieves state-of-the-art performance when compared to existing specialized and multi-task models across multiple tasks.
@@ 2 Related Work
Facial analysis tasks: Facial analysis tasks involve face parsing [28, 98, 11, 86, 122, 62], landmarks detection [46, 37, 41, 125, 55], head pose estimation [124, 90, 115, 12], facial attributes recognition [64, 57, 80, 123], age/gender/race estimation [8, 36, 39, 43] and landmarks visibility prediction [54, 35]. These tasks hold significance in various applications such as face swapping [13, 60], face editing [127], de-occlusion [111], 3D face reconstruction [102], driver assistance [59], human-robot interaction [81], retail [1], face verification and identification [84, 85],image generation [109], image retrieval [113] and surveillance [22, 61]. Specialized models excel in their respective tasks but cannot be easily integrated with other tasks due to the need for extensive task-specific pre-processing [47, 125]. Generally, these models under-perform when applied to tasks beyond their specialization as their design is specific to their designated tasks. Some works [119, 56, 121, 26] perform multiple tasks simultaneously but utilize the additional tasks for guidance or auxiliary loss calculation to enhance the performance of the primary task. HyperFace [70] and AllinOne [71] are early models that explore multi-task learning with the aim of performing multiple tasks. HyperFace utilizes multi-scale features from different layers of CNNs, and is capable of landmarks detection, head pose estimation and gender estimation, while AllinOne additionally performs face recognition and age estimation. Recent models, such as SwinFace [66], QFace [83], and Faceptor [67], leverage transformer models with learnable tokens to perform multi-task learning. However, they perform fewer tasks, neglecting more complex ones such as head pose estimation, landmark estimation, and face parsing. The proposed FaceXFormer addresses all these tasks, alongside others, and achieves state-of-the-art performance in majority of them.
@@ 3 FaceXFormer
Figure 2: Overview of our proposed framework. The FaceXFormer employs an encoder-decoder architecture, extracting multi-scale features from the input face image $\mathbf{I}$, and fusing them into a unified representation $\mathbf{F}$ via MLP-Fusion. Task tokens $\mathbf{T}$ are processed alongside face representation $\mathbf{F}$ in the FaceX Decoder $\mathbf{FXDec}$, resulting in refined task-specific tokens $\mathbf{\hat{T}}$. These refined tokens are then used for task-specific predictions by passing through the unified head.
In our framework, we follow a standard encoder-decoder structure as illustrated in Fig. 2. For an input face image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, we extract coarse to fine-grained multi-scale features $\mathbf{S}_{i}$, where $i$ belongs to the $i$-th encoder output. To learn a unified face representation $\mathbf{F}$, these multi-scale features are then fused using a lightweight MLP-Fusion $\mathbf{M}$ module. Following fusion, we initialize a series of task-specific tokens $\mathbf{T}=\langle T_{1},\dots,T_{n}\rangle$, with each $t_{i}$ representing a face task. Afterward, we initialize task tokens $\mathbf{T}=\langle T_{1},\dots,T_{n}\rangle$, where $T_{i}$ denotes each task. Face tokens $\mathbf{F}$ and task tokens $\mathbf{T}$ are then processed by a Parameter-efficient Decoder $\mathbf{FXDec}$ where task tokens are attended with face tokens to learn relevant task representation.
@@ 3.1 Multi-scale Encoder
Lightweight MLP-Fusion: Assigning each feature-map $\mathbf{S}_{i}$ to each face task is sub-optimal; rather, learning a unified face representation is more optimal and parameter-efficient. Following [107], we utilize a Lightweight MLP-Fusion module $\mathbf{M}$ to generate a fused face representation from the multi-scale features $\{\mathbf{S}_{i}\}_{i=1}^{n}$. In this framework, each feature map $\mathbf{S}_{i}$ is initially passed through a separate MLP layer, standardizing the channel dimensions across scales to facilitate fusion. The transformed features are then concatenated and passed through a fusion MLP layer to aggregate a fused representation $\mathbf{F}$ as follows:
@@
where $D_{i}$ and $D_{t}$ are the multi-scale feature channel dimensions of $\mathbf{S}_{i}$ and the target channel dimension, respectively. The lightweight MLP-fusion design ensures minimal computational overhead while maintaining the ability to perform efficient feature fusion, which is crucial for real-time application based face analysis tasks.
@@ 3.2 FaceX Decoder
Detection transformer (DETR) [10] employs object tokens to learn bounding box predictions for each object. Inspired by this approach, we introduce Task Tokens, whereby each task token is designed to learn specific facial tasks leveraging the fused face representation. However, existing decoders such as DETR [10] and Deformable-DETR [129] are computationally intensive, impacting runtime significantly. To address this, we propose a FaceX ($\mathbf{FXDec}$) a parameter-efficient decoder designed to efficiently model the task tokens with face tokens. Specifically, each task token learns a task-related representation by interacting with other task tokens $\mathbf{T}$ and face tokens $\mathbf{F}$, enhancing the overall representation. The Parameter-Efficient Decoder comprises three main components: 1) Task Self-Attention, 2) Task-to-Face Cross-Attention, and 3) Face-to-Task Cross-Attention as illustrated in Fig. 2.
@@
Task-to-Face Cross-Attention (TFCA): The Task-to-Face Cross-Attention module allows each task token to interact with the fused face representation $\mathbf{F}$. This enables each task token to gather information relevant to its specific facial task from thed fuse face features. In this module, the fused face representation $\mathbf{F}$ acts as both key and value, while the task tokens serve as queries. The updated task token $\hat{T}_{i}$ is then computed as follows:
@@ 3.3 Unified-Head
In Unified-Head, the task tokens are processed to obtain corresponding task predictions. As shown in Fig. 2, the output face tokens $\mathbf{\hat{F}}$ and task tokens $\mathbf{\hat{T}}$ are processed through a Task-to-Face Cross-Attention mechanism to obtain final refined features. Then, the output tokens are fed into their corresponding task heads. The task head for landmark detection is a hourglass network and for head pose estimation is a regression MLP, while the tasks of estimating age, gender, race, expression, visibility, and attributes prediction utilize classification MLPs. For face parsing, we leverage the output $\mathbf{\hat{F}}$ and process it through an upsampling layer, then perform a cross-product with the face parsing token to obtain a segmentation map. The number of tokens for segmentation corresponds to the total number of classes. For landmark prediction, it corresponds to the number of landmarks (i.e., $68$). For head pose estimation, the number of tokens is $9$, representing the $3\times 3$ rotation matrix. For other tasks, one token is used for each.
@@ 3.4 Multi-Task Training
$L=\lambda_{seg}L_{seg}$ $+\lambda_{lnd}L_{lnd}+\lambda_{hpe}L_{hpe}+\lambda_{attr}L_{attr}$
$+\lambda_{a}L_{a}+\lambda_{g/r}L_{g/r}+\lambda_{exp}L_{exp}+\lambda_{vis}L_{vis}$
where $L_{seg}$ is the mean of dice loss [82] and Cross-Entropy (CE) loss for face parsing, $L_{lnd}$ is STAR loss [125] for landmarks prediction, $L_{hpe}$ is geodesic loss [115] for head pose estimation, $L_{g/r}$ is CE loss for gender/race estimation, $L_{a}$ is mean of L1 loss and CE loss for age estimation, $L_{exp}$ is CE loss for facial expression recognition, and $L_{attr}$ and $L_{vis}$ are Binary Cross-Entropy with logits loss for attributes prediction and face visibility prediction respectively.
@@
Table 2: Performance comparison for face parsing on the CelebAMask-HQ dataset [38]. The symbol $\times$ indicates that the model does not perform the corresponding task. Red = First Best, Blue = Second Best.

### v3

@@ Abstract
In this work, we introduce FaceXFormer, an end-to-end unified transformer model capable of performing ten facial analysis tasks within a single framework. These tasks include face parsing, landmark detection, head pose estimation, attribute prediction, age, gender, and race estimation, facial expression recognition, face recognition, and face visibility. Traditional face analysis approaches rely on task-specific architectures and pre-processing techniques, limiting scalability and integration. In contrast, FaceXFormer employs a transformer-based encoder-decoder architecture, where each task is represented as a learnable token, enabling seamless multi-task processing within a unified model. To enhance efficiency, we introduce FaceX, a lightweight decoder with a novel bi-directional cross-attention mechanism, which jointly processes face and task tokens to learn robust and generalized facial representations. We train FaceXFormer on ten diverse face perception datasets and evaluate it against both specialized and multi-task models across multiple benchmarks, demonstrating state-of-the-art or competitive performance. Additionally, we analyze the impact of various components of FaceXFormer on performance, assess real-world robustness in “in-the-wild” settings, and conduct a computational performance evaluation. To the best of our knowledge, FaceXFormer is the first model capable of handling ten facial analysis tasks while maintaining real-time performance at $33.21$ FPS.
@@ 1 Introduction
Face analysis is a crucial problem as it has broad range of application such as face verification and identification [92, 93], surveillance [25], face swapping [14], face editing [137], de-occlusion [120], 3D face reconstruction [111], retail [1], image generation [118] and face retrieval [122]. Facial analysis tasks (Figure 1 include face parsing [33, 107], landmarks detection [51, 135], head pose estimation [134, 13], facial attributes recognition [71, 63], age/gender/race estimation [9, 48], facial expression recognition [85], face recognition [39], and face visibility prediction [60, 41]. Therefore, developing a generalized and robust face model for all tasks is a crucial and longstanding problem in the face community.
Figure 1: FaceXFormer an end-to-end unified transformer model for 10 different facial analysis tasks such as face parsing, landmark detection, head pose estimation, attributes recognition, age, gender, and race estimation, facial expression recognition, face recognition, and face visibility prediction.
Why Unified Model ? In recent years, significant advancements have been made in facial analysis, developing state-of-the-art methods and face libraries for various tasks  [134, 135, 48, 13, 120, 14]. Despite these methods achieving promising performance, they cannot be integrated into a single pipeline due to their specialized model designs and task-specific pre-processing techniques. Furthermore, deploying multiple specialized models simultaneously is computationally intensive and impractical for real-time applications, leading to increased system complexity and resource consumption. These challenges emphasis the need for a unified model that can concurrently handle multiple facial analysis tasks efficiently (see Table 1). A single model capable of addressing multiple facial tasks is desirable because it: (1) learns a robust and generalized face representation capable of handling “in-the-wild” images; (2) intra-task modeling helps the models to learn task-invariant representation; and (3) simplifies deployment pipelines by reducing computational overhead and enabling faster inference.
Methods FP LD HPE Attr Age Gen Race Vis Exp FR
@@
Table 1: Comparison with representative methods under different task settings. Our proposed FaceXFormer can perform various facial analysis tasks in a single model. FP - Face Parsing, LD - Landmarks Detection, HPE - Head Pose Estimation, Attr - Attributes Recognition, Age - Age, Gen - Gender, Race - Race Estimation, Exp - Facial Expression Recognition, FR - Face Recognition, and Vis - Face Visibility Prediction
Proposed FaceXFormer Architecture: To this end, we introduce FaceXFormer, an end-to-end unified model designed for ten different facial analysis tasks, as depicted in Figure 1. These tasks include face parsing, landmark detection, head pose estimation, attributes recognition, age/gender/race estimation, facial expression recognition, face recognition and face visibility prediction. FaceXFormer enables task unification by leveraging transformers and learnable tokens as its core components. Specifically, we employ a transformer-based encoder-decoder structure, where the encoder extracts hierarchical face representations and fuses them using a MLP fusion module. The fused features are then processed in the decoder, where each facial analysis task is represented by a unique learnable token, allowing for the simultaneous and effective processing of multiple tasks. In particular, we propose a lightweight decoder, FaceX, which processes both face and task tokens together using bi-directional cross-attention mechanism (Section 3.2), enabling the model to learn robust face representations that generalize across various tasks. The bi-directional cross-attention mechanism enables a 2-layer lightweight decoder, allowing the model to operate in real time. After modeling the intra-task and face-token relationships in the FaceX decoder, the task tokens are fed into a unified head, which converts these task tokens into corresponding task predictions.
Our extensive experiments demonstrate that FaceXFormer achieves state-of-the-art or competitive performance compared to specialized models and existing multi-task models across multiple benchmarks, while supporting more tasks than any previous multi-task model. Moreover, we show that our model effectively handles images “in the wild”, demonstrating its robustness and generalizability across ten different tasks. This robustness is critical for real-world applications where uncontrolled conditions and diverse inputs are common. FaceXFormer achieves state-of-the-art performance at $33.21$ FPS, representing a significant $69.44$% speed boost over prior multi-task models, making it highly suitable for real-world applications.
@@
    We introduce FaceXFormer, a unified transformer-based framework capable of simultaneously processing ten different facial analysis tasks, achieving real-time performance of $33.21$ FPS.
@@
    We propose FaceX, a lightweight decoder that employs the proposed bi-directional cross-attention mechanism, enabling joint processing of face and task tokens.
@@
    We conduct extensive experiments and analyses to demonstrate that our approach achieves state-of-the-art performance with reduced inference time compared to specialized and multi-task models across multiple tasks.
@@ 2 Related Work
Facial analysis tasks: Facial analysis tasks involve face parsing [33, 12, 130, 69], landmarks detection [51, 135, 61], head pose estimation [134, 98, 124, 13], facial attributes recognition [71, 63, 88, 133], age/gender/race estimation [9, 42, 45, 48], facial expression recognition [47, 46], face recognition [100, 16] and face visibility prediction [60, 41]. These tasks hold significance in various applications such as face swapping [14, 67], face editing [137], de-occlusion [120], 3D face reconstruction [111], driver assistance [66], human-robot interaction [89], retail [1], face verification and identification [92, 93], image generation [118], image retrieval [122] and surveillance [25, 68]. Specialized models excel in their respective tasks but cannot be easily integrated with other tasks due to the need for extensive task-specific pre-processing [52, 135]. Generally, these models under-perform when applied to tasks beyond their specialization as their design is specific to their designated tasks. Some works [127, 62, 129, 30] perform multiple tasks simultaneously but utilize the additional tasks for guidance or auxiliary loss calculation to enhance the performance of the primary task.
Multi-task learning for face analysis: HyperFace [77] and AllinOne [78] are early convolution-based models that aim to perform multiple tasks. Recent multi-task frameworks, such as QFace [91] and Faceptor [74], are also inspired from DETR [11] and propose a unified model structure consisting of learnable tokens. However, these previous works differ from the proposed method in several key aspects, as summarized in Table 2. Specifically, QFace [91] employs a feature fusion module that uses stage embeddings to aggregate features from the encoder. Faceptor [74] introduces a layer-attention mechanism to fuse features from different encoder layers and incorporates two separate decoders. Both methods, employ a 9-layer transformer decoder, also Faceptor additionally includes a Pixel Decoder. These architectural components increase computational overhead, resulting in slower inference times. In contrast, FaceXFormer proposes a bi-directional cross-attention mechanism, which enables efficient task-specific feature extraction from face tokens resulting in a 2-layer lightweight decoder. This design choice is the primary reason for FaceXFormer’s superior speed and performance. Notably, unlike previous methods, FaceXFormer does not rely on face-specific pertaining backbone.
QFace [91] Faceptor [74] FaceXFormer
Tasks 4 7 10
Model Size - 178.9 M 109.29 M
Face Pretrained Yes Yes No
FPS (fp32) - 14.30 33.21
Decoder Query2Label [54] Pixel Decoder & FaceX
Transformer Decoder
Decoder Layers 9 9 2
Table 2: Comparison of multi-task face analysis methods.
@@ 3 FaceXFormer
Figure 2: Overview of our proposed framework. The FaceXFormer employs an encoder-decoder architecture, extracting multi-scale features from the input face image $\mathbf{I}$, and fusing them into a unified representation $\mathbf{F}$ via MLP-Fusion. Task tokens $\mathbf{T}$ are processed alongside face representation $\mathbf{F}$ in the FaceX Decoder $\mathbf{FXDec}$, resulting in refined task-specific tokens $\mathbf{\hat{T}}$. These refined tokens are then used for task-specific predictions by passing through the unified head. FaceXFormer performs ten tasks, including face parsing, landmark detection, head pose estimation, attribute prediction, age, gender, and race estimation, facial expression recognition, face recognition, and face visibility prediction, achieving state-of-the-art performance at a real-time FPS of $33.21$.
In our framework, we follow a standard encoder-decoder structure as illustrated in Fig. 2. For an input face image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, we extract coarse to fine-grained multi-scale features $\mathbf{S}_{i}$, where $i$ belongs to the $i$-th encoder output. To learn a unified face representation $\mathbf{F}$, these multi-scale features are then fused using a MLP-Fusion $\mathbf{M}$ module. Following fusion, we initialize a series of task-specific tokens $\mathbf{T}=\langle T_{1},\dots,T_{n}\rangle$, with each $t_{i}$ representing a face task. Afterward, we initialize task tokens $\mathbf{T}=\langle T_{1},\dots,T_{n}\rangle$, where $T_{i}$ denotes each task. Face tokens $\mathbf{F}$ and task tokens $\mathbf{T}$ are then processed by a lightweight Decoder $\mathbf{FXDec}$ where task tokens are attended with face tokens to learn relevant task representation.
@@ 3.1 Multi-scale Encoder
MLP-Fusion: Assigning each feature-map $\mathbf{S}_{i}$ to each face task is sub-optimal; rather, learning a unified face representation is more optimal and parameter-efficient. Following [116], we utilize a MLP-Fusion module $\mathbf{M}$ to generate a fused face representation from the multi-scale features $\{\mathbf{S}_{i}\}_{i=1}^{n}$. In this framework, each feature map $\mathbf{S}_{i}$ is initially passed through a separate MLP layer, standardizing the channel dimensions across scales to facilitate fusion. The transformed features are then concatenated and passed through a fusion MLP layer to aggregate a fused representation $\mathbf{F}$ as follows:
@@
where $D_{i}$ and $D_{t}$ are the multi-scale feature channel dimensions of $\mathbf{S}_{i}$ and the target channel dimension, respectively. The MLP-fusion design ensures minimal computational overhead (983k parameters) while maintaining the ability to perform efficient feature fusion, which is crucial for real-time application based face analysis tasks.
@@ 3.2 FaceX Decoder
Detection transformer (DETR) [11] employs object tokens to learn bounding box predictions for each object. Inspired by this approach, we introduce Task Tokens, whereby each task token is designed to learn specific facial tasks leveraging the fused face representation. However, existing decoders such as DETR [11] and Deformable-DETR [139] are computationally intensive, impacting runtime significantly. To address this, we propose FaceX ($\mathbf{FXDec}$) a lightweight decoder designed to efficiently model the task tokens with face tokens. Specifically, each task token learns a task-related representation by interacting with other task tokens $\mathbf{T}$ and face tokens $\mathbf{F}$, enhancing the overall representation. The Lightweight Decoder comprises of three main components: 1) Task Self-Attention, 2) Task-to-Face Cross-Attention, and 3) Face-to-Task Cross-Attention as illustrated in Figure 2.
@@
Task-to-Face Cross-Attention (TFCA): The Task-to-Face Cross-Attention module allows each task token to interact with the fused face representation $\mathbf{F}$. This enables each task token to gather information relevant to its specific facial task from the fused face features. In this module, the fused face representation $\mathbf{F}$ acts as both key and value, while the task tokens serve as queries. The updated task token $\hat{T}_{i}$ is then computed as follows:
@@ 3.3 Unified-Head
In Unified-Head, the task tokens are processed to obtain corresponding task predictions. As shown in Figure 2, the output face tokens $\mathbf{\hat{F}}$ and task tokens $\mathbf{\hat{T}}$ are processed through a Task-to-Face Cross-Attention mechanism to obtain final refined features. Then, the output tokens are fed into their corresponding task heads. The task head for landmark detection is a hourglass network, for head pose estimation is a regression MLP, and for face recognition is PartialFC [4], while the tasks of age, gender and race estimation, facial expression recognition, face visibility prediction, and attributes prediction utilize classification MLPs. For face parsing, we leverage the output $\mathbf{\hat{F}}$ and process it through an upsampling layer, then perform a cross-product with the face parsing token to obtain a segmentation map. The number of tokens for segmentation corresponds to the total number of classes. For landmark prediction, it corresponds to the number of landmarks (i.e., $68$). For head pose estimation, the number of tokens is $9$, representing the $3\times 3$ rotation matrix. For other tasks, one token is used for each.
@@ 3.4 Multi-Task Training
[…]
where $L_{seg}$ is the mean of dice loss [90] and Cross-Entropy (CE) loss for face parsing, $L_{lnd}$ is STAR loss [135] for landmarks prediction, $L_{hpe}$ is geodesic loss [124] for head pose estimation, $L_{g/r}$ is CE loss for gender/race estimation, $L_{a}$ is mean of L1 loss and CE loss for age estimation, $L_{exp}$ is CE loss for facial expression recognition, $L_{fr}$ is ArcFace [16] loss for face recognition, and $L_{attr}$ and $L_{vis}$ are Binary Cross-Entropy with logits loss for attributes prediction and face visibility prediction respectively.
@@
Table 3: Performance comparison for face parsing on the CelebAMask-HQ dataset [44]. The symbol $\times$ indicates that the model does not perform the corresponding task. Red = First Best, Blue = Second Best. $\times$ indicates a model that doesn’t perform the task.